[Lerftest] LERF Status

Kevin Jordan jordan at jlab.org
Thu Sep 3 16:51:06 EDT 2020

The gun knife switch was shut off this morning per procedure & Bubba placed his lock on it, the crew chief lock is in the process of being moved from the gun valve to the power supply. Carlos & Bubba will do a re-cesiation starting ~7:00 AM Friday.

The vault will be left in Controlled Access overnight.

All required PSS certifications are now complete, many thanks to the RF group for getting the zones in HV, and to DC power group for completing the 1X power supply checkout & certification & of course to the SSG!!!

Shukui will check out the drive laser Friday AM as well.

Once Carlos is finished in the vault we will go to Power Permit and try to establish gradient in all of the cavities; Buncher, Quarter, & Zone 2. This work will cease around noon when Mike Drury is ready to locally retune the LCLS-II cavities for upcoming microphonics studies. I checked all of the zone 2 tuners this morning and all function properly - the "exercise script" must be killed (note was sent to Pam). The vault will be dropped to Restricted Access for this activity & left that way through the weekend.

The first 4 of 8 Solid State Amplifiers (SSA) have been staged for removal next week.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/lerftest/attachments/20200903/98d42aef/attachment.html>

More information about the Lerftest mailing list